A stone projected horizontally from the top of a tower with a speed of 5m/s. It lands on the ground level of the horizontal distance of 20m from the top of the tower. Calculate the height of the tower. g= 10m/s²

1
Expert's answer
2022-01-21T13:48:45-0500

Given:

"v_0=5\:\rm m/s"

"l=20\:\rm m"

"g=10\:\rm m/s^2"


The height of a tower

"H=\frac{gt^2}{2}"

Time of falling

"t=l\/v_0"

So

"H=\frac{gl^2}{2v_0^2}=\frac{10*20^2}{2*5^2}=80\:\rm m"